# The Virtual Reality Koinos Method: Analyzing Virtual Reality Collaboration from the perspective of communication models

### v1

@@ Abstract
Understanding which factors could influence co-presence in Virtual Reality could help develop more qualitative social interactions, or social interactions that generate similar sensations, emotions and feelings than the ones generated during Face-to-Face interactions. Co-presence is studied since the beginning of Virtual Reality (VR); though, no consensus is identified on what factors could influence it, except the consensus on the definition of “being there together” inside the Virtual Environment. In this paper, we introduce the Koinos method to explain social interactions in VR through communication models, (i) theoretically, and (ii) on two VR experiments that change the virtual partner social and physical representations. These analyses lead us to propose an equation to predict and help manage the sense of co-presence in VR.
@@
Virtual Reality , Social interactions , Communication , Methodology , Collaboration , Avatars
@@
The use of communication models helps to better understand social interactions dynamics in Virtual Reality (VR).
This work proposes the Koinos method: a method adapting communication models with the specificities of VR in the aim to analyze the dynamics of social interactions in VR.
The analyze of two collaborative VR experiment with the Koinos method indicates that avatar appearance does not seems prior on co-presence level and task completion. In addition, prejudices and ideas associate to the virtual collaborator seems to guide the behavior and the perceived quality of social interactions.
@@ 1 Introduction
Among the Computer-Mediated Communication (CMC) devices, Virtual Reality (VR) can be seen as a medium that allows to reproduce the most faithfully social interactions qualia, comparing to the Face-to-Face (FtF) interactions. A quale (singular of qualia) is a set of emotions, sensations and feelings, created and updating from different experiences. Therefore, the qualia of the FtF interactions are continuously updating, social interactions being central to our daily life. Since our first days, according to Meltzoff & Moore [1], the interactions with others allow us to construct ourselves. In VR, to evaluate how much social interactions are similar in VR and in FtF, a common measure used is the co-presence level. Defined as the feeling of “being there together” in the Virtual Environment (VE) (Gallagher & Frith [2]), understanding which factors could influence it is a primary subject to allow more qualitative social interactions in VR. However, the literature has not shown clear results on this subject, regarding the complexity and the cross-disciplinary fields involved.
@@
The study of communication models, such as the social interaction dynamics with Newcomb’s ABX model, Newcomb [3], and the Cooperative Model of Human Communication of Tomasello [4], could help better understand what could influence co-presence in VR.
@@
In this theoretical study, we introduce the Koinos method by first applying these communication models on VR situations, and then analyze two VR experiments that use the same collaborative task but with different partner entities and appearances, with the communication models. These analyses lead us to propose a VR co-presence equation that could help predict and manage better the feeling of co-presence, and therefore could help create more qualitative social interaction VR applications.
2 Related work on Communication Models
@@
Social interactions involve inter-personal communication, that takes place in a social context. Indeed, a conversation will not be conducted in the same way depending on the place it happens for instance, or the subject of a meeting. In addition, to transmit a message, both verbal and non-verbal communications enable smooth social interaction, allowing to understand each other, and to take into account the internal states of others. The ABX model of Newcomb [3] represents these complex social interactions, by representing both the relationships between the individuals, and the social context (Figure 1).
@@
Figure 1: Newcomb’s ABX Model [3]. A and B can represent individuals and X represents a message transmitted into a context. The double arrow between A and B represents the reciprocal influence, the verbal and non-verbal communication as well as the relationship between them.
@@
A and B can represent two individuals, the double arrow represents the reciprocal influence between the two, the verbal and non-verbal communication, as well as the relationship between them. X can represent the message transmitted into a social context. The theory of Newcomb [3] indicates that a certain balance exists in this triangle; and that a change in one element changes the entire social interaction dynamics, leading to the emergence of a new equilibrium. For example, we can see A and B as coworkers, exchanging about a work subject. If A has a promotion and becomes the chief of B, the relationship between A and B will change; therefore, the conversation about the same work subject will change too. This model can also be used by considering A or/and B as a group, and an other dynamics will occur than when considering A and B as individuals.
@@
The Cooperative Model of Human Communication of Tomasello [4] brings insights and complementary information about the inter-personal interaction. This theory takes into account the individual goals and values in the communication act. To achieve these goals, individuals may need the help, exchange of information, or sharing attitudes with others, which is the social intention. The social intention is then communicated, through the communication intention, and allows to share common goals. Therefore, the use of references from the external world can be used to communicate the message and the intention, which is the referential intention. Based on norms of cooperation, that indicate how to cooperate with others, the other will try to understand why a communication is attempted, the cooperative reasoning will help understand that. Therefore, to understand the referential intention element, an individual will base his/her reasoning on the common ground, the elements both individuals share. Finally, the individual will decide if he/she wants to cooperate or not.
@@
Tomasello proposed that all human beings use this process in communication. Therefore, one factor that could influence the understanding in communication, can be the common ground. The common ground corresponds to elements that interactors have in common. Tomasello [4] describes three levels to evaluate if the common ground is strong or not:
@@
    The common ground is based on immediate perceptual environment or on a shared experience from the past;
@@
Tomasello [4] explains that the common ground is stronger when based on a shared previous experience, in the case of a top-down process and when it is based on implicit knowledge. Therefore, the “strength” of the common ground depends of the relationship between interactors, on the situation, and on the common knowledge.
@@
In addition, Tomasello [4] indicates that, the more the persons share a strong common ground, the less the use of communication elements will be necessary to understand and communicate with the other. We can cite the case of a wife and a husband, twins, or best friends for instance, who can understand each other in just a glance. Therefore, we can suppose that to understand a communication intention of someone with a very low common ground, more communication will be necessary, both verbal and non-verbal (e.g., asking a glass of water in a plane to a waiter from another country).
@@
Previous works have used the common ground perspective to bring new insights to the understanding of social interactions dynamics in social VE. Axelsson et al. [5] compared symmetric and asymmetric hardware setups (i.e., a CAVE system and computers) on the communication between participants. By comparing the results with the different setups, the authors indicate four factors that allow to better understand the others in a collaborative task: hearing the partner, seeing the partner and partner’s interact objects, sharing the view of the environment with the partner and understanding how the partner interacts with the VE. This study supports Tomasello’s theory [4], that indicates that if individuals do not know each other (i.e., low common ground), more complex communication features are necessary to understand the intention of the others.
@@
It is possible to create a common ground with someone who is unknown in a VE, at the condition that an affordance is present between the task objectives and the interactions inside the VE, according to Shami et al. [6].
@@
In addition, the model of Roos et al. [7] proposes a model to represent the social dynamics with mediated communication. This model explains that the behavior of users is constrained by three elements: Personal Possibilities, Technical Possibilities and Social Possibilities; and results in a Situated Social Structure. The authors explain that the inter-personal interaction is inherently social and collaborative, and that the smoothness of their interaction is used as evidence to update their beliefs about the relationship with others. Consequently, disruption issued from technical issues, could lead to social expectation disruption about the partner’s interaction, which arises questions about the common goals and identity as a collective. This model puts into perspective social context with technical and personal ones.
@@
Finally, for social interactions in VR, the feeling of “being in the VE together”, Schroeder [8], also named co-presence, can be related to social interactions. Indeed, the Cooperative Model of Human Communication of Tomasello [4] can only occur if each individual considers the other as a partner for collaboration. Therefore, if a virtual avatar does not provide the feeling that it is really controlled by a human, it could be challenging to experience collaboration or social interaction. Therefore, Schroeder presents a model to represent the continuum of connected presence, that could be applied on various devices allowing communication. A distinction is made between co-presence and connected presence, by the shared environment that connected presence has. This model considers co-presence as the feeling of being with others through mediated communication devices in which users do not share the same environment (i.e., phone call); and use the term of connected presence when the environment is shared (i.e., social VE and computer). This distinction is made in the case of asymmetric collaboration. Our study concerning social interactions in VR, participants will use symmetrical mediated device. Therefore, we will use the term of co-presence for the feeling of being in the VE together, as described by Biocca et al. [9].
3 Methodology: the Koinos method – applying communication models in VR-mediated systems
@@
To better understand the social interactions dynamics occurring in VR, we suggest to apply the Newcomb’s ABX model. X will correspond to the message transmitted, that can be related to a task, inside a VE, through the VR hardware. The relationship between A and B will be therefore mediated; the communication between the interactors will take place through their respective avatars. However, regarding the cooperative model of human communication, a link still exists between A and B: the common ground. Both communication models are combining in a single model depicted in Figure 2, that describe the social interaction dynamics in VR.
@@
Figure 2: VR social interaction dynamics viewed through communication models: Newcomb’s ABX model [3] and Cooperative model of Human Communication [4]
@@
This model, allows to understand that, even if the communication is mediated by avatars, a link is still present between users A and B, through the common ground. Based on past literature, we can propose that this common ground could be of different “strengths” that can influence the communication smoothness and the social interaction.
@@
Resulting from Tomasello’s theory [4] , we propose a grid that allows to estimate the common ground “strength” (Table 1).
Only one possibility on the two choices in each part:
@@
Tomasello [4] explains that the common ground will be stronger when individual share experience from the past, when the situation is based on a top-down process, and when they shared implicit common knowledge. The values of “1” or “2” were chosen according to these results. For instance, a value of “2” is related to a shared experience from the past, and a value of “1” is accorded to the immediate perceptual environment. This allows to evaluate the strength of the common ground of a given situation and individuals, and could change with time. The “0” value is set in the cases where the chosen common ground item does not actually exist.
According to the proposed grid, the highest value for the common ground is 6, when the common ground is based on shared experience from the past, when the situation implements a top-down process and when the common knowledge is based on implicit common knowledge. For instance, co-workers, who work together from several years on a common project, and know what the other knows about this project, induce the maximum value of common ground. Conversely, the lowest value can be 0; therefore, no common ground is present. A low value can be identified, for instance if strangers walk in a street, and hear a strange noise, it could induce a value of 3 (one in each 3 categories), or 2 if the strangers are tourists from different countries for instance (explicit common cultural knowledge is then rated as 0).
[…]
3.1 Toward a prediction of the co-presence level?
From past research and our proposed Koinos Method, we question if the social interactions dynamics in VR could be predicted, especially the co-presence level. Indeed, by considering that co-presence reflects the perceived quality of the social interactions, the fact that with a high common ground, individuals do not need complex communication, and that the social interactions happen in a constraint VE, we wonder if identifying and studying each of these factors could help predict the co-presence level. The social interactions allowed and implemented in the VE seems to have an impact in this prediction, thus we will use the term of communication cues to define the communication means that are included in the VR application, that could be verbal and non-verbal, such as voice, text chat, eyes or face tracking and animation.
@@ 3.1 Toward a prediction of the co-presence level?
4 Applying the Koinos Method on VR experiments that vary social representations and the avatar appearance of a virtual partner in a collaborative task
@@ 4.1 VR experiment 1: Effect of the social representation on co-presence and behavior
A first study was designed to explore whether it is possible to induce different co-presence levels with the social representation theory, based on Abric [10] and Poinsot et al. [11]. We have explored the following research questions: Will co-presence be affected by different social representations of the virtual collaborator? Will participants collaborate in the same way with a human collaborator or with an AI-based one? Will user’s behavior be correlated to the co-presence level they experience?
@@ 4.1.1 Experimental design
The main experimental condition was to induce different levels of co-presence, in order to study what other factors could be related to it. Participants experienced two VR sessions, one with a Human virtual partner and the other with an AI-controlled avatar, in random order. In real, in both cases, participants completed the experiment together without knowing it. They were separated in different rooms without the possibility to see or hear each other.
A collaborative task was developed in VR for Oculus Quest 2, in which the goal was to gain points by completing patterns in collaboration. Each participant had a specific task: either following a pattern with the hand, or reproducing the movement of the collaborator (Figure 3). In order to generate social representations of each entity, the avatar appearance was congruent to the entity avatar represented, i.e., a human shape in the “Human” condition (Figure 4, left and right), and an avatar representing an AI for the “AI” condition (Figure 4, center). Previous to the experiment, an online survey was conducted in order to identify which visual representation is associated to AI, and what are the words or sentences associated to AI. A total of 105 answers was analyzed, allowing to identify that the visual representation of AI is a mix of the algorithm idea with a human representation, such as mathematical formula with a humanoid robot that look like thinking. The social representation of AI results of terms such as “algorithms”, “help” or “non-human” terms.

### v2

# The Virtual Reality Koinos Method: Analysis of Symmetrical Dyadic Collaboration in Virtual Reality from the perspective of communication models

@@ Abstract
Understanding which factors influence co-presence in Virtual Reality can help develop more qualitative social interactions, or social interactions that experienced similar sensations, emotions, and feelings to those generated during Face-to-Face interactions. Co-presence has been studied since the beginning of Virtual Reality (VR); however, no consensus has been established on the factors that influence it, except the consensus on the definition of ”being there together” within the Virtual Environment. In this paper, we introduce the Koinos method to explain social interactions in VR through communication models, (i) theoretically, and (ii) on two VR symmetrical dyadic collaborative experiments that change the virtual partner’s social and physical representations. These analyses lead us to propose an equation that predict and helps manage the sense of co-presence in VR.
@@
Virtual Reality , Social interactions , Communication , Methodology , Co-presence , Avatars
@@
The use of communication models helps to better understand social interaction dynamics in Virtual Reality (VR).
This work proposes the Koinos method: a method adapting communication models with the specificities of VR, with the aim of analyzing the dynamics of social interactions in VR.
The analysis of two symmetrical dyadic collaborative VR experiments with the Koinos method indicates that avatar appearance does not affect the co-presence level and task completion.
Social representations associated with the virtual collaborator guide the behavior and the perceived quality of social interactions, resulting in significantly lower co-presence with an AI-controlled avatar than with a Human-controlled avatar.
@@ 1 Introduction
Among the Computer-Mediated Communication (CMC) devices, Virtual Reality (VR) can be seen as a medium that allows for reproducing social interactions qualia, compared to the Face-to-Face (FtF) interactions. A quale (singular of qualia) is a set of emotions, sensations, and feelings, created and updated from different experiences. Therefore, the qualia of the FtF interactions are continuously updating, and social interactions are central to our daily life. Since our first days, according to Meltzoff & Moore [meltzoff1977imitation], the interactions with others allow us to construct ourselves. In VR, to evaluate how much social interactions are similar in VR and FtF, a common measure used is the co-presence level. Defined as the feeling of “being there together” in the Virtual Environment (VE) (Gallagher & Frith [gallagher2003functional]), understanding which factors could influence it is a primary subject to allow more qualitative social interactions in VR. However, the literature has not provided precise results on this subject, due to the complexity and the cross-disciplinary fields involved.
@@
The study of communication models, such as the social interaction dynamics with Newcomb’s ABX model, Newcomb [newcomb1953approach], and the Cooperative Model of Human Communication of Tomasello [tom2008], could help better understand what could influence co-presence in VR.
@@
In this theoretical study, we introduce the Koinos method by first applying these communication models to VR situations, and then analyzing two VR experiments that use the same collaborative task but with different partner entities and appearances, using the communication models. These analyses lead us to propose a VR co-presence equation that could help predict and manage better the feeling of co-presence, and therefore could help create more qualitative social interaction VR applications.
2 Related work
@@
Social interactions involve interpersonal communication, which occurs within a social context. Indeed, a conversation will not be conducted in the same way depending on the place it happens, for instance, or the subject of a meeting. In addition, to transmit a message, both verbal and non-verbal communications enable smooth social interaction, allowing us to understand each other, and to take into account the internal states of others. The ABX model of Newcomb [newcomb1953approach] represents these complex social interactions by representing both the relationships between the individuals and the social context (Figure 1).
@@
Figure 1: Newcomb’s ABX Model [newcomb1953approach]. A and B can represent individuals and X represents a message transmitted into a context. The double arrow between A and B represents the reciprocal influence, the verbal and non-verbal communication as well as the relationship between them.
@@
A and B can represent two individuals; the double arrow represents the reciprocal influence between the two, encompassing bot verbal and non-verbal communication, as well as the relationship between them. X can represent the topic discussed, such as a third person, or any subject. This model shows that A and B have a ”simultaneous orientation” to each other and to X. The theory of Newcomb [newcomb1953approach] indicates that a certain balance exists in this triangle; and that a change in one element changes the entire social interaction dynamics, leading to the emergence of a new equilibrium. For example, we can see A and B as coworkers, exchanging about a work subject. If A receives a promotion and becomes the chief of B, the relationship between A and B will change; therefore, the conversation about the same work subject will also change.
@@
The Cooperative Model of Human Communication by Tomasello [tom2008] offers insights and complementary information about the interpersonal interaction. This theory considers the individual goals and values in the communication act. To achieve these goals, individuals may need the help, exchange of information, or sharing of attitudes with others, which is the social intention. The social intention is then communicated through the communication intention, allowing for the sharing of common goals. Therefore, the use of references from the external world can be used to communicate the message and the intention, which is the referential intention. Based on norms of cooperation, which indicate how to cooperate with others, the other will try to understand why communication is attempted, and the cooperative reasoning will help understand that. Therefore, to understand the referential intention element, an individual will base his/her reasoning on the common ground, the elements both individuals share. Finally, the individual will decide whether to cooperate or not.
@@
Tomasello proposed that all human beings use this process in communication. Therefore, one factor that influences the understanding in communication is the common ground. The common ground corresponds to elements that interactors have in common. Tomasello [tom2008] describes three levels to evaluate if the common ground is strong or not:
@@
    The common ground is based on the immediate perceptual environment or on a shared experience from the past;
@@
Tomasello [tom2008] explains that the common ground is firmer when based on a shared previous experience, in the case of a top-down process, and when it is based on implicit knowledge. Therefore, the “strength” of the common ground depends on the relationship between interactors, on the situation, and on the common knowledge.
@@
In addition, Tomasello [tom2008] indicates that the more the persons share a strong common ground, the less the use of communication elements will be necessary to understand and communicate with one another. We can cite the case of a wife and a husband, twins, or best friends, for instance, who can understand each other at just a glance. Therefore, to understand the communication intention of someone with a very low common ground, more communication will be necessary, both verbal and non-verbal (e.g., asking for a glass of water in a plane to a server from another country).
@@
Previous works have used the common ground perspective to bring new insights to the understanding of social interaction dynamics in social VE. Axelsson et al. [axelsson2003communication] compared symmetric and asymmetric hardware setups (i.e., a CAVE system and computers) on the communication between participants. By comparing the results with the different setups, the authors indicate four factors that allow for better understanding of the others in a collaborative task: hearing the partner, seeing the partner and the partner’s interacted objects, sharing the view of the environment with the partner, and understanding how the partner interacts with the VE. This study supports Tomasello’s theory [tom2008], which indicates that when individuals do not know each other (i.e., low common ground), more complex communication features are necessary to understand the intentions of others.
@@
It is possible to establish common ground with someone who is unknown in a VE, provided an affordance exists between the task objectives and the interactions within the VE, according to Shami et al. [shami2011places].
@@
In addition, the model proposed by Roos et al. [roos2024social] aims to represent the social dynamics through mediated communication. This model explains that the behavior of users is constrained by three elements: Personal Possibilities, Technical Possibilities, and Social Possibilities, and results in a Situated Social Structure. The authors explain that interpersonal interaction is inherently social and collaborative, and that the smoothness of their interaction is used as evidence to update their beliefs about the relationship with others. Consequently, disruptions caused by technical issues could lead to social expectations being disrupted about the partner’s interaction, which raises questions about common goals and identity as a collective. This model puts the social context into perspective alongside technical and personal ones.
@@
Finally, for social interactions in VR, the feeling of “being in the VE together” (also known as co-presence) in Schroeder study [schroeder2006being], can be related to social interactions. Indeed, the Cooperative Model of Human Communication of Tomasello [tom2008] can only occur if each individual considers the other as a partner for collaboration. Therefore, if a virtual avatar does not provide the feeling that it is really controlled by a human, it could be challenging to experience collaboration or social interaction. Therefore, Schroeder presents a model to represent the continuum of connected presence, which can be applied to various devices that allow communication. A distinction is made between co-presence and connected presence based on the shared environment that characterizes connected presence. This model considers co-presence as the feeling of being with others through mediated communication devices in which users do not share the same environment (i.e., phone call); and uses the term connected presence when the environment is shared (i.e., social VE and computer). This distinction is made in the case of asymmetric collaboration. In our study concerning social interactions in VR, participants will use a symmetrical mediated device. Therefore, we will use the term ”co-presence” for the feeling of being in the VE together, as described by Biocca et al. [biocca2003toward].
Therefore, the co-presence level, measured with a questionnaire [biocca2003toward], can be seen as an indicator of the perceived quality of the social interaction. Co-presence represents the feeling that avatars are the representations of their human users, that the avatars are the humans. According to the Theory of Mind [gallagher2003functional] and previous communication models presented above, a qualitative social interaction only occurs when we consider the other as human. In other words, participants in VR perceive social interaction in VR as equivalent in quality to face-to-face interaction. Therefore, to study co-presence in VR, it is necessary to conduct social interactions, which can be seen as a reciprocal influence in a situated context [stebe2007risques]. A meta-analysis found that social influence with avatars or agents only emerges in collaborative or competitive tasks, and does not occur in neutral ones [fox2015avatars]. In both collaboration and competition, the simultaneous orientation of the actors is directed towards the same object of the collaboration or the competition. Although co-presence can be measured in any situation or task, collaboration and competition tasks allow us to control the simultaneous orientation of actors on the same subject and on each other, thereby reproducing Newcomb’s model [newcomb1953approach].
3 Koinos method: applying communication models in VR-mediated systems
@@
To better understand the social interaction dynamics occurring in VR, we suggest applying Newcomb’s ABX model. X will correspond to the message transmitted, which can be related to a task within a VE, through the VR hardware. The relationship between A and B will therefore be mediated; the communication between the interactors will take place through their respective avatars. However, regarding the cooperative model of human communication, a link still exists between A and B: the common ground. Both communication models are combined in a single model depicted in Figure 2, which describes the social interaction dynamics in VR, with a focus on the embodied avatar.
@@
Figure 2: VR social interaction dynamics viewed through communication models: Newcomb’s ABX model [newcomb1953approach] and Cooperative model of Human Communication [tom2008].
@@
This model enables us to understand that, even if communication is mediated by avatars, a link still exists between users A and B through the common ground. Based on past literature, we can propose that this common ground may have different “strengths” that can influence the communication smoothness and the social interaction.
@@
Resulting from Tomasello’s theory [tom2008], we propose a grid that allows for estimating the common ground “strength” (Table 1).
Only one possibility in each of the three columns:
@@
Tomasello [tom2008] explains that the common ground will be stronger when individuals share experience from the past, when the situation is based on a top-down process, and when they share implicit common knowledge. The values of “1” or “2” were chosen according to these results. For instance, a value of “2” is associated with a shared experience from the past, and a value of “1” is assigned to the immediate perceptual environment. The grid allows for the evaluation of the strength of the common ground in a given situation and individuals, which can change over time. The “0” value is set in the cases where the chosen common ground item does not exist.
According to the proposed grid, the highest value for the common ground is 6, when the common ground is based on shared experience from the past, when the situation implements a top-down process, and when the common knowledge is based on implicit common knowledge. For instance, co-workers who work together for several years on a common project, and know what the other knows about this project, induce the maximum value of common ground. Conversely, the lowest value can be 0; therefore, no common ground is present. A low value can be identified, for instance if strangers walk in a street, and hear a strange noise, it could induce a value of 3 (one in each three categories), or two if the strangers are tourists from different countries for instance (explicit common cultural knowledge is then rated as 0).
The proposed methodology aims to identify the various factors influencing social interaction dynamics in a VR application (Figure 2). The use of the common ground evaluation grid (Table 1) allows for the estimation of the strength of the common ground, and a better understanding of the link between participants and the social interaction dynamics that occur. We propose to name this methodology the Koinos Method. The term Koinos (ϰοινος) is derived from ancient Greek, signifying “common”, “shared” or “collective”. In the context of this methodology, Koinos represents the underlying principles of shared understanding, collaborative interaction, and collective experience; elements that are central to the dynamics of social interaction in VR.
Toward a prediction of the co-presence level?
From past research and our proposed Koinos Method, we question if the social interaction dynamics in VR could be predicted, especially the co-presence level. Indeed, by considering that co-presence reflects the perceived quality of the social interactions, the fact that with a high common ground, individuals do not need complex communication, and that the social interactions happen in a constrained VE, we wonder if identifying and studying each of these factors could help predict the co-presence level. The social interactions allowed and implemented in the VE seems to have an impact in this prediction, thus we will use the term of communication cues to define the communication means that are included in the VR application, that could be verbal and non-verbal, such as voice, text chat, eyes or face tracking and animation.
@@
4 Applying the Koinos Method on collaborative VR experiments
@@ 4.1 VR experiment 1: Effect of the social representation on co-presence and behavior
A first study was designed to explore whether it is possible to induce different co-presence levels with the social representation theory, based on Abric [abric1989etude] and Poinsot et al. [poinsot2022effect], and an online survey on the social representation of AI. We have explored the following research questions: Will co-presence be affected by different social representations of the virtual collaborator? Will participants collaborate in the same way with a Human collaborator or with an AI-based one? Will user’s behavior be correlated to the co-presence level they experience?
@@ 4.1.1 Experimental design
The main experimental condition was to induce different levels of co-presence, in order to study what other factors could be related to it. Participants experienced two VR sessions, one with a Human-controlled avatar and the other with an AI-controlled avatar, in random order. In reality, in both cases, participants completed the experiment together without knowing it, being the Human or AI partner of each other without knowing it. They were separated in different rooms without the possibility of seeing, talking, or hearing each other.
A collaborative task was developed in VR for Oculus Quest 2, in which the goal was to gain points by completing patterns in collaboration, taking into account the speed and accuracy. Each participant had a specific task: either following a pattern with the hand or reproducing the movement of the collaborator (Figure 3). In order to generate social representations of each entity, the avatar appearance was congruent to the entity avatar represented, i.e., a human shape for the Human-controlled avatar (Figure 4, left and right), and an avatar representing an AI for the AI-controlled avatar (Figure 4, center). Facial expressions were not included and participants could communicate with gestures between two completions of the pattern.
Previous to the experiment, an online survey was conducted in order to identify which visual representation is associated with AI, and what words or sentences are associated with AI. A total of 105 answers were analyzed, allowing us to identify that the visual representation of AI is a mix of the algorithm idea with a human representation, such as a mathematical formula with a humanoid robot that looks like it is thinking. The social representation of AI results in terms such as “algorithms”, “help”, or “non-human” terms.